# Asynchronous Microphone Array Calibration using Hybrid TDOA Information

### 12 Mar 2024 v2

@@ Abstract
Asynchronous Microphone array calibration is a prerequisite for most audition robot applications. In practice, the calibration requires estimating microphone positions, time offsets, clock drift rates, and sound event locations simultaneously. The existing method proposed Graph-based Simultaneous Localisation and Mapping (Graph-SLAM) utilizing common TDOA, time difference of arrival between two microphones (TDOA-M), and odometry measurement, however, it heavily depends on the initial value. In this paper, we propose a novel TDOA, time difference of arrival between adjacent sound events (TDOA-S), combine it with TDOA-M, called hybrid TDOA, and add odometry measurement to construct Graph-SLAM and use the Gauss-Newton (GN) method to solve. TDOA-S is simple and efficient because it eliminates time offset without generating new variables. Simulation and real-world experiment results consistently show that our method is independent of microphone number, insensitive to initial values, and has better calibration accuracy and stability under various TDOA noises. In addition, the simulation result demonstrates that our method has a lower Cramér-Rao lower bound (CRLB) for microphone parameters, which explains the advantages of my method.
@@ I INTRODUCTION
Microphone arrays can equip robots with sound source localization and tracking abilities, etc [19, 1, 2]. A prerequisite for realizing the above functionalities is to accurately calibrate the array geometric information [3]. A common approach to the above calibration problem is to utilize the time difference of arrival measurements between microphone pairs (TDOA-M) from a series of sound events. Earlier methods require the clock synchronization of all microphones [4, 5]. To overcome the limitation, recent studies, including [6, 7, 8], have estimated microphone positions with an asynchronous factor: time offsets.
Figure 1: Scenario difference between TDOA-M (a) and TDOA-S (b). (a) shows sound source localization using TDOA-M and known microphone locations, and (b) shows microphone localization using TDOA-S and known sound event locations obtained from odometry measurements.
During calibration, one can obtain the relative position measurements between adjacent sound events from the odometer onboard the robot (which acts as a moving sound source) and use them to improve the calibration accuracy. Following the above idea, based on TDOA-M and odometry measurements, an extended Kalman filter-based simultaneous localization and mapping (EKF-SLAM) method has been proposed in [13] to estimate microphone positions, time offsets, and sound source positions simultaneously. However, the impact of the other asynchronous factor, clock drift rates, has not been considered. In [9, 10, 11], a batch SLAM-based method [12] has been presented to also estimate the clock drift rate and this method requires a good initial value to achieve accurate calibration results, which is impractical.
Motivated by sound source localization utilizing TDOA-M, we obtain the time difference of arrival between adjacent sound events (TDOA-S) to locate microphones. Fig. 1 illustrates the difference between TDOA-M and TDOA-S in the calibration scene. Our main contributions are stated as follows.
@@
    A novel and efficient measurement: TDOA-S without time offset is proposed and we also introduce a simple method of extracting TDOA-S in practice. To our best knowledge, this is the first time TDOA-S has been proposed in the literature and used in calibrating robot audition systems. The idea can be generalized to other sensing modalities.
@@
    Based on hybrid TDOA information (which combines TDOA-S and TDOA-M) and odometry measurements, we have proposed a batch SLAM-based method to jointly estimate the asynchronous microphone array parameters (including microphone positions, time offsets, clock drift rates) and sound source positions.
  • $\bullet$
    We have designed simulations and real-world experiments to validate that our method is independent of microphone number, less sensitive to initialization, has higher accuracy and stability under various TDOA noises, and has lower CRLB for microphone parameters. We further open-source our code and real-world data to benefit the community [^1] https://github.com/Chen-Jacker/Hybrid-TDOA-Calib.git.
@@ II The proposed METHOD
Assume there are $N$ microphones. Denote the $i$-th microphone location, time offset, and clock drift rate as $\mathbf{x}_{i}$, $\tau_{i}$, and $\delta_{i}$ respectively. The unknown microphone parameters are
@@
where $\tau_{i,1}=\tau_{i}-\tau_{1}$, $i>1$. There are $K$ sound events and the $j$-th sound event location and emitting time are $\mathbf{s}_{j}$ and $t_{j}$, respectively. Without loss of generality, in our method, the coordinate frame is established by sound event positions, called $Sound$ frame, $\mathbf{s}_{1}=\mathbf{0},(\mathbf{s}_{2})_{y}=(\mathbf{s}_{2})_{z}=(\mathbf{s}_{3})_{z}=0$. Sound source parameters that need to be estimated are
@@ II-A1 Derivation
TDOA-S is derived from the time of arrival (TOA) model that considers two asynchronous parameters: time offset and clock drift rate in microphones. In the absence of noise, the arrival time detected by $i$-th microphone for the $j$-th sound event, $T_{i,j}$ is shown below,
@@
where $c$ is the sound speed, $\tau_{i}$ and $\delta_{i}$ represent the shift and scaling of the temporal frame of $i$-th microphone with respect to (w.r.t.) the absolute temporal frame, respectively. The former is caused by different startup moments in different microphones, and the latter is caused by the sampling rate mismatch between the microphone’s actual and absolute sampling rates [15], which can be modeled as a scale constant between a microphone temporal frame and the absolute temporal frame. If the mismatch does not exist for $i$-th microphone ($\delta_{i}=0$), the TOA model is the same as the common TOA that only considers time offset [3].
In indoor calibration scenarios, the distance between the microphone and sound events does not generally exceed 10 meters. In most cases, the clock drift rate and time offset are less than $10^{-4}$ and 0.1s respectively. Therefore, the nonlinear term $\delta_{i}(\frac{||\mathbf{x}_{i}-\mathbf{s}_{j}||}{c}+\tau_{i})$ is so small that can be ignored. After this simplification,
@@
Therefore, TDOA-S, the time difference reaching a microphone between adjacent sound events, is expressed as $T^{S}_{i,j}=\tilde{T}_{i,j+1}-\tilde{T}_{i,j}$ and $\Delta t_{j}=t_{j+1}-t_{j}$. The measurement model of $T^{S}_{i,j}$ is
@@
where $j<K$ and $\Delta t_{j}$ is known.
@@ II-A2 Extraction
There are two steps in obtaining TDOA-S, each visualized in Fig. 2. Initially, all audio segments containing a calibration signal in a single-channel microphone are captured for a time window. The rough time delay of adjacent calibration signals ($T_{rough}$) equals to difference between the left endpoint of the adjacent windows containing the calibration signal (Fig. 2a). Next, align the adjacent windows and perform GCC-PHAT [14] to obtain the precise delay ($T_{pre}$) (see Fig. 2b). Finally, combines the rough delay and precise delay to obtain the overall delay ($T_{rough}+T_{pre}$), which is TDOA-S and equal to the difference between two consecutive moments of arrival.
Figure 2: Visualization of acquiring the rough delay: $T_{rough}$ (a) and precise delay: $T_{pre}$ (b). The red/green box represent the capture window obtaining the current/next recorded calibration signal.
@@ II-B Calibration using Hybrid TDOA
II-B1 Hybrid TDOA Measurementss
@@
where $\tau_{i,1}=\tau_{i}-\tau_{1}$, $\delta_{i,1}=\delta_{i}-\delta_{1}$ ($i>1$) and $t_{j}=t_{j}-t_{1}$ as assume $t_{1}=0$ without loss of generality. $t_{j}$ is known because the sound emitting time interval is known. The TDOA-M formula 6 is equivalent to the TDOA formula in [11]. Hence, without noise, the total hybrid TDOA measurements are
@@
Under Gaussian noise $\mathbf{v}_{j}\sim N(\mathbf{0},\sigma^{2}_{odo}\mathbf{I_{3}})$, the odometry measurements are $\mathbf{m}=[\mathbf{m}_{1},\mathbf{m}_{2},...\mathbf{m}_{K-1}]^{T}$ with $\mathbf{m}_{j}$ being defined as follows
@@
where $j<K$.
@@ II-B2 Nonlinear Least Squares solved by GN method
From the perspective of batch SLAM, nodes are the locations of a series of sound events (robot pose without orientation) and microphone array (landmark) with positions and asynchronous parameters, while edges are odometry measurements and hybrid TDOA measurements. Because any microphone observes every sound event, data association is easily achieved. One can then construct the corresponding nonlinear least squares based on maximum likelihood estimate (MLE) and then use the Gauss-Newton (GN) method to estimate microphone array positions, time offsets, clock drift rates, and the sound event locations simultaneously. To be specific, define parameters $\mathbf{x}=[\mathbf{x}_{mic},\mathbf{s}]^{T}$, measurement $\mathbf{z}=[\mathbf{t}^{H},\mathbf{m}]^{T}$ and measurement function $\mathbf{f}(\mathbf{x})=[\mathbf{T}^{H},\Delta\mathbf{s}]^{T}$. The minimum of the nonlinear least squares shown below,
@@
where $\mathbf{W}=diag(\sigma^{2}_{tdoa}\mathbf{I}_{N(K-1)+K(N-1)},\sigma^{2}_{odo}\mathbf{I}_{3K-3})$. The GN method is a gradient-based iterative solution method to solve the nonlinear least squares problem and outputs $\mathbf{x}$ following (1) and (2). For performing source localization tasks after calibration, we need to convert $\mathbf{x}_{mic}$ following (1) to $\mathbf{x}_{mic}$ following (17). The details of the transformation are shown in Appendix A.
@@ II-C Computation of CRLB
CRLB is a popular and powerful tool for analyzing parameter estimation errors, as it provides a lower bound on the estimated parameter variance for any unbiased estimator. In this paper, we compute the CRLB of our method and the method [10] and determine which TDOA makes its method have the lower CRLB for microphone parameters. For nonrandom vector parameters, the CRLB states that the covariance matrix of an unbiased estimator is bounded as follows [16],
@@
In our method, we consider the GN solver for the nonlinear least squares (10) as the unbiased estimator and the CRLB matrix of $\mathbf{x}_{mic}$ following (1), called $\mathbf{x}^{S}_{mic}$ here, is defined as $\mathbf{C}_{\mathbf{x}_{mic}^{S}}=\mathbf{C}(1:5N-1,1:5N-1)$, which is the submatrix of $\mathbf{C}$ w.r.t. $\mathbf{x}$. Then, we need to obtain CRLB for $\mathbf{x}_{mic}$ following (17), called $\mathbf{x}^{M}_{mic}$. The affine transformation between $\mathbf{x}^{M}_{mic}$ and $\mathbf{x}^{S}_{mic}$ is represented below
@@
In $\mathbf{C}_{\mathbf{x}_{mic}^{M}}$, we extract diagonal elements corresponding to the CRLB for $\mathbf{x}_{mic}^{M}$. Then we define an indicator $D_{CRLB}$ to evaluate and $D_{CRLB_{i}}$ can be represented as CRLB for $i$-th microphone location, offset, or clock drift rates.
$D_{CRLB}=\sqrt{\frac{\sum_{i=2}^{N}\ D_{CRLB_{i}}}{N-1}}.$ (16)
@@ III SIMULATIONS
We next present simulations to validate the advantages of our method: independence of microphone number (Part A), less insensitivity to initial values (Part B), better calibration accuracy and stability under various TDOA noises (Part C), and lower CRLB for microphone parameters (Part D). For comparative analysis, we use the existing calibration method using TDOA-M in 3D version [10].
@@ III-1 Setup
We design two motion trajectories of a sound source. One has the space of 3m×$\times$3m×$\times$3m with eight sound events (trajectory 1) and the other has the space of 2m×$\times$6m×$\times$2m with 10 sound events (trajectory 2).
@@
In “True $\mathbf{x}_{mic}$”, “random” means microphone locations are randomly generated in the corresponding trajectory space and $|\tau_{i,1}|\leq 0.1s$, $|\delta_{i}|\leq 10^{-4}s$. In “Initial $\mathbf{x}_{mic}$”, “random” means both microphone and sound event locations are randomly generated in the corresponding trajectory space and asynchronous parameters set to be zero. $\sigma_{init}$ are standard deviations of zero-mean Gaussian noises adding into the true positions as the initial values of both microphone and sound event locations. In trajectory 1, $\sigma_{init}=0m,1m,2m,3m$ and in trajectory 2 $\sigma_{init}=0m,2m,4m,6m$. Simulation under different numbers of microphones (Part A), various initial value noises (Part B), and several TDOA noises (Part C/D) repeat 200 times in each trajectory and the results of two trajectories are combined to analyze.
III-2 Metric
The average root mean square of estimated microphone location errors (Loc. err.), time offset errors (Off. err.) and clock drift rates errors (Dri. err.) are evaluated in the $Mic.$ frame whose definition is in Appendix A.
Figure 3: Box plot of estimation errors of microphone parameters in simulations: microphone locations (a),(d),(g), time offsets (b),(e),(h), and clock drift rates (c),(f),(i) under various microphone numbers, initial values noises, and TDOA noises respectively.
III-A Simulation Results
There is an observation in Fig. 3a-c: as the number of microphones changes, the calibration performance for microphone parameters of our method remains basically unchanged. However, the performance of [10] shows significant changes and approaches that of our method as the number increases.
Fig. 3d-f shows that the estimation performance for microphone parameters of our method remains unchanged under different initial value noises. However, microphone parameters estimated by [10] exhibit an increase in estimation error as the initial values noise increases.
In Fig. 3g-i, we can observe that our method has better accuracy and stability in estimating microphone parameters under three levels of TDOA noises, as we have lower median and IQR values for each box.
Table I confirms that our method estimates the CRLB for microphone parameters to be smaller under a variety of TDOA noises.
TABLE II: CRLB RESULTS UNDER VARIOUS TDOA NOISES

### 19 Mar 2024 v3

@@ Abstract
Asynchronous microphone array calibration is a prerequisite for most audition robot applications. A popular solution to the above calibration problem is the batch form of Simultaneous Localisation and Mapping (SLAM), using the time difference of arrival measurements between two microphones (TDOA-M), and the robot (which serves as a moving sound source during calibration) odometry information. In this paper, we introduce a new form of measurement for microphone array calibration, i.e. the time difference of arrival between adjacent sound events (TDOA-S) with respect to the microphone channels. We propose to combine TDOA-S and TDOA-M, called hybrid TDOA, together with odometry measurements for bath SLAM-based calibration of asynchronous microphone arrays. Simulation and real-world experiment results consistently show that our method is more independent of microphone number, less sensitive to initial values (when using off-the-shelf algorithms such as Gauss-Newton iterations), and has better calibration accuracy and robustness under various TDOA noises. In addition, the simulation result demonstrates that our method has a lower Cramér-Rao lower bound (CRLB) for microphone parameters. To benefit the community, we open-source our code and data at https://github.com/zcj808/Hybrid-TDOA-Calib.
@@ I INTRODUCTION
Microphone arrays can equip robots with sound source localization and tracking abilities, etc [1, 2, 3, 4, 5, 6]. A prerequisite for realizing the above functionalities is to calibrate the array geometric information accurately [7]. A common approach to the above calibration problem is to utilize the time difference of arrival between microphone pairs (TDOA-M) from a series of sound events to estimate both microphone and sound source locations. Earlier methods require the clock synchronization of all microphones [8, 9]. To overcome the limitation, recent studies, including [10, 11, 12], take into account the initial time offset between microphone channels.
Figure 1: Differences between TDOA-M (a) and TDOA-S (b).
During calibration, one can obtain the relative position measurements between adjacent sound events from the odometer onboard the robot (which acts as a moving sound source) and use them to improve the calibration accuracy. Following the above idea, based on TDOA-M and odometry measurements, an extended Kalman filter-based simultaneous localization and mapping (EKF-SLAM) method is proposed in [13] to estimate microphone positions, time offsets, and sound source positions simultaneously. However, the impact of the other asynchronous factor, i.e. clock drift rate, is not considered. In [14, 15, 16], a batch SLAM-based method [17] is presented to jointly estimate microphone locations, time offsets, clock drift rates, and sound event positions. However, when using off-the-shelf algorithms such as Gauss-Newton iterations, this method requires a good initial value to achieve reasonable calibration results.
In this paper, we introduce a new form of measurement for microphone array calibration, i.e. the time difference of arrival between adjacent sound events (TDOA-S) with respect to the microphone channels. Fig. 1 shows the difference between TDOA-M and TDOA-S in the calibration scene. We propose to combine TDOA-S with TDOA-M, called hybrid TDOA, with odometry measurements for microphone array calibration. Our main contributions are stated as follows.
@@
    We introduce TDOA-S and present a simple method of extracting TDOA-S in practice. To our best knowledge, this is the first time TDOA-S has been proposed in the literature and used in calibrating robot audition systems. The idea can be generalized to other sensing modalities.
@@
    We propose a batch SLAM-based calibration method utilizing hybrid TDOA information and odometry measurements to jointly estimate the asynchronous microphone array parameters (microphone positions, time offsets, clock drift rates) and sound source positions. Extensive simulations and real-world experiments have been conducted to validate the advantages of the proposed method. The results show that compared to [17], our method is more independent of microphone number, less sensitive to initialization, has higher accuracy and robustness under various TDOA noises, and has lower CRLB for microphone parameter estimates.
@@ II THE PROPOSED METHOD
Assume there are $N$ microphones. Denote the $i$-th microphone location, time offset, and clock drift rate as $\mathbf{x}_{i}$, $\tau_{i}$, and $\delta_{i}$ respectively. There are $K$ sound events and the $j$-th sound event location is $\mathbf{s}_{j}$. Without loss of generality, the coordinate frame is established by sound event positions, called $Sound$ frame, $\mathbf{s}_{1}=\mathbf{0},(\mathbf{s}_{2})_{y}=(\mathbf{s}_{2})_{z}=(\mathbf{s}_{3})_{z}=0$. The unknown microphone parameters in the $Sound$ frame are
@@
where $\tau_{i,1}=\tau_{i}-\tau_{1}$, $i>1$. Sound source parameters that need to be estimated are
@@ II-A1 Derivation
TDOA-S is derived from the time of arrival (TOA) model that considers two asynchronous parameters: time offset and clock drift rate in microphones. In the absence of noise, the arrival time detected by $i$-th microphone for the $j$-th sound event, $T_{i,j}$ is shown below
@@
where $c$ is the sound speed and $t_{j}$ is the emitting time of $j$-th sound event. $\tau_{i}$ and $\delta_{i}$ represent the shift and scaling of the temporal frame of $i$-th microphone with respect to (w.r.t.) the absolute temporal frame, respectively. The former is caused by different startup moments in different microphones. The latter is caused by the sampling rate mismatch between the microphone’s actual and absolute sampling rates [18], which can be modeled as a scale constant between a microphone’s temporal frame and the absolute temporal frame. If the mismatch does not exist for $i$-th microphone ($\delta_{i}=0$), the TOA model is the same as the common TOA that only considers time offset [7].
In indoor calibration scenarios, the distance between the microphone and sound events does not generally exceed 10 meters. In most cases (Table I in [16]), the clock drift rate and time offset are less than $10^{-4}$ and 0.1s respectively. Therefore, the term $\delta_{i}(\frac{||\mathbf{x}_{i}-\mathbf{s}_{j}||}{c}+\tau_{i})$ is negligible and can be ignored. After this simplification, $T_{i,j}$ becomes
@@
Therefore, TDOA-S is expressed as $T^{S}_{i,j}=\tilde{T}_{i,j+1}-\tilde{T}_{i,j}$. The measurement model of $T^{S}_{i,j}$ is
@@
where $j<K$ and $\Delta t_{j}=t_{j+1}-t_{j}$ can be obtained accurately since the speaker installed on the robot is controllable.
@@ II-A2 Extraction
There are two steps in obtaining TDOA-S, each visualized in Fig. 2. Initially, all audio segments containing a calibration signal in a single-channel microphone are captured for a time window. The rough time delay of adjacent calibration signals ($T_{rough}$) is equal to the difference between the left endpoint of the adjacent windows containing the calibration signal (Fig. 2a). Next, we align the adjacent windows and perform GCC-PHAT [19] to obtain the precise delay ($T_{pre}$) (see Fig. 2b). Finally, one combines the rough delay and precise delay to obtain the overall delay ($T_{rough}+T_{pre}$), which is equal to the difference between two consecutive moments of arrivals.
Figure 2: Visualization of acquiring the rough delay: $T_{rough}$ (a) and precise delay: $T_{pre}$ (b). The red/green box represents the capture window obtaining the current/next recorded calibration signal.
@@ II-B Calibration using Hybrid TDOA
II-B1 Hybrid TDOA Measurements
@@
where $\tau_{i,1}=\tau_{i}-\tau_{1}$, $\delta_{i,1}=\delta_{i}-\delta_{1}$ ($i>1$). Assume $t_{1}=0$ without loss of generality, for $j>1$, $t_{j}=t_{j}-t_{1}=\sum^{j}_{k=2}\Delta t_{k-1}$. The TDOA-M formula (6) is equivalent to the TDOA formula in [16]. Hence, without noise, the total hybrid TDOA measurements are
@@
Under Gaussian noise $\mathbf{v}_{j}\sim N(\mathbf{0},\sigma^{2}_{odo}\mathbf{I_{3}})$, the odometry measurements are $\mathbf{m}=[\mathbf{m}_{1},\mathbf{m}_{2},...,\mathbf{m}_{K-1}]^{T}$ with $\mathbf{m}_{j}$ being defined as follows
@@
where $j<K$. Denote $\Delta\mathbf{s}=[\Delta\mathbf{s}_{1},\Delta\mathbf{s}_{2},...,\Delta\mathbf{s}_{K-1}]^{T}$.
@@ II-B2 Nonlinear Least Squares solved by GN method
From the perspective of batch SLAM, nodes are the locations of a series of sound events (robot pose without orientation) and microphone array (landmark) with positions and asynchronous parameters, while edges are odometry measurements and hybrid TDOA measurements. Note that during the calibration process, any microphone observes every sound event. One can then construct the corresponding nonlinear least squares based on maximum likelihood estimate (MLE) and then use the Gauss-Newton (GN) method to estimate microphone array positions, time offsets, clock drift rates, and the sound event locations simultaneously. Denote the unknown parameters $\mathbf{x}=[\mathbf{x}_{mic},\mathbf{s}]^{T}$, measurements $\mathbf{z}=[\mathbf{t}^{H},\mathbf{m}]^{T}$ and measurement function $\mathbf{f}(\mathbf{x})=[\mathbf{T}^{H},\Delta\mathbf{s}]^{T}$. The minimum of the nonlinear least squares (LS) is shown below
@@
where $\mathbf{W}=diag(\sigma^{2}_{tdoa}\mathbf{I}_{N(K-1)+K(N-1)},\sigma^{2}_{odo}\mathbf{I}_{3K-3})$. For solving the above optimization problem, the GN method is usually used. Moreover, for performing source localization tasks after calibration, we need to convert $\mathbf{x}_{mic}$ in $Sound$ frame to $\mathbf{x}_{mic}$ in $Mic.$ frame. The details of the transformation are shown in Appendix A.
@@ II-C Computation of CRLB
CRLB is a popular and powerful tool for analyzing parameter estimation errors, as it provides a lower bound on the estimated parameter variance for any unbiased estimator. For nonrandom vector parameters, the CRLB states that the covariance matrix of an unbiased estimator is bounded as follows [20],
@@
In our method, we consider the GN solver for the nonlinear least squares (10) as the unbiased estimator and the CRLB matrix of $\mathbf{x}_{mic}$ in $Sound$ frame, called $\mathbf{x}^{S}_{mic}$ here, is defined as $\mathbf{C}_{\mathbf{x}_{mic}^{S}}=\mathbf{C}(1:5N-1,1:5N-1)$, which is the submatrix of $\mathbf{C}$ w.r.t. $\mathbf{x}$. Then, we need to obtain CRLB for $\mathbf{x}_{mic}$ in $Mic.$ frame, called $\mathbf{x}^{M}_{mic}$. The affine transformation between $\mathbf{x}^{M}_{mic}$ and $\mathbf{x}^{S}_{mic}$ is represented below
@@
In $\mathbf{C}_{\mathbf{x}_{mic}^{M}}$, we extract diagonal elements corresponding to the CRLB for $\mathbf{x}_{mic}^{M}$. Then we define an indicator $D_{CRLB}$
$D_{CRLB}=\sqrt{\frac{\sum_{i=2}^{N}\ D_{CRLB_{i}}}{N-1}},$ (16)
where $D_{CRLB_{i}}$ is represented as CRLB for $i$-th microphone location, offset, or clock drift rate in $Mic.$ frame.
@@ III SIMULATIONS
We next present simulations to validate the advantages of our method: independence of the number of microphones (Part A), less insensitivity to initial values (Part B), better calibration accuracy and robustness under various TDOA noises (Part C), and lower CRLB for microphone parameters (Part D). For comparative analysis, we use the calibration method [15] using TDOA-M in 3D.
@@ III-1 Setup
We design three motion trajectories of a sound source. The first one has the space of 3m×$\times$3m×$\times$3m with 8 sound events (trajectory 1), the second has the space of 2m×$\times$6m×$\times$2m with 10 sound events (trajectory 2), and the third one possesses the space of 4m×$\times$4m×$\times$2m with 14 sound events (trajectory 3).
@@
In “True $\mathbf{x}_{mic}$”, “random” means microphone locations are randomly generated in the corresponding trajectory space and $|\tau_{i,1}|\leq 0.1s$, $|\delta_{i}|\leq 10^{-4}s$. In “Initial $\mathbf{x}_{mic}$”, “random” means both microphone and sound event locations are randomly generated in the corresponding trajectory space, and asynchronous parameters are set to be zero. $\sigma_{init}$ are standard deviations (SDs) of zero-mean Gaussian noises added into the true positions as the initial values of both microphone and sound event locations. In trajectory 1, $\sigma_{init}=0m,1m,2m,3m$ and in trajectory both 2 and 3, $\sigma_{init}=0m,2m,4m,6m$. Simulation under different numbers of microphones (Part A), various initial value noises (Part B), and several TDOA noises (Part C/D) repeat 200 times in each trajectory and the results of three trajectories are combined to analyze.
III-2 Evaluation Metric
The average root mean square errors of the estimated microphone locations (Loc. err.), time offset (Off. err.), and clock drift rates (Dri. err.) are evaluated in the $Mic.$ frame, whose definition is in Appendix A.
Figure 3: Box plot of estimation errors for microphone parameters in simulations: microphone locations (a),(d),(g), time offsets (b),(e),(h), and clock drift rates (c),(f),(i) under four microphone numbers, i.e. 4, 6, 8, and 10, four initial values noise SDs under three trajectories, i.e. 0m/0m, 1m/2m, 2m/4m, and 3m/6m, and three TDOA noise SDs, i.e. $5\times 10^{-5}s$, $10^{-4}s$, and $5\times 10^{-4}s$, respectively. Here, “$x$m/$y$m” means combining the estimation result of trajectory 1 under $\sigma_{init}=x$m and results of both trajectory 2 and 3 under $\sigma_{init}=y$m.
III-3 Results
It can be observed from Fig. 3a-c that as the number of microphones changes, the calibration performance for microphone parameters of our method remains basically unchanged. However, the performance of [15] shows significant changes and approaches that of our method as the number increases. Fig. 3d-f shows that the estimation performance for microphone parameters of our method remains unchanged under different initial value noises. However, microphone parameters estimated by [15] exhibit an increase in estimation error as the initial values noise increases. In Fig. 3g-i, we can observe that our method has better accuracy and robustness in estimating microphone parameters under three levels of TDOA noises, as we have lower median and IQR values for each box. Table I confirms that our method estimates the CRLB for microphone parameters to be smaller under various TDOA noises.
TABLE II: CRLB results under various TDOA noises (Bold means better)